I'm homeschooling my cousin.
I don't know where to start with this problem: 2(c-3)+5 = 3c-3
email: chrisen_c@yahoo.com

Answer by dabanfield@earthlink.net(8)

1.) 2(c-3)+5 = 3c-3

I assume you are solving for c.
Expanding the left side of equation 1.) we have
2.) 2c - 6 + 5 = 3c - 3
Simplify 2.):
3.) 2c - 1 = 3c - 3
Subtract 2c from both sides of 3.):
4.) 2c - 2c - 1 = 3c - 2c - 3
Simplify 4.):
5.) -1 = c - 3
Add 3 to both sides of equation 5.)
-1 + 3 = c -3 + 3
2 = c